OpenStudy (anonymous):

Use the binomial theorem to write the binomial expansion of (c-4)^5

OpenStudy (anonymous):

The binomial theorem for expansion follows: \[(x+y)^{n}=\left(\begin{matrix}n \\ 0\end{matrix}\right)x^ny^0 + \left(\begin{matrix}n \\ 1\end{matrix}\right)x^{n-1}y^1+\left(\begin{matrix}n \\ 2\end{matrix}\right)x^{n-2}y^2+...+\left(\begin{matrix}n \\ {n-1}\end{matrix}\right)x^1y^{n-1}+\left(\begin{matrix}n \\ n\end{matrix}\right)x^0y^n\]

OpenStudy (anonymous):

Do you know how to expand from that?

OpenStudy (anonymous):

yup. thanks!

OpenStudy (anonymous):

part of it was cut off at the end, should be \[...+\left(\begin{matrix}n \\ {n-1}\end{matrix}\right)x^1y^{n-1} + \left(\begin{matrix}n \\ n\end{matrix}\right)x^0y^n\]

OpenStudy (anonymous):

that's ok. I figured that out. thanks again